# An Embedding-based Approach to Inconsistency-tolerant Reasoning with Inconsistent Ontologies

### v2

@@ 3. Preliminaries
Definition 2 (scoring function). A scoring function s associates with an ontology $\mathrm{\Sigma}$ and an axiom $\alpha\in\mathrm{\Sigma}$ a non-negative real number $s(\mathrm{\Sigma},\alpha)$ which is equal to 00 if and only if $\alpha$ is trivial (i.e., such that $\alpha$ is a tautology or a contradiction).
@@
#$\#mc(\mathrm{\Sigma},\alpha)=\left|\left\{\mathrm{\Sigma}_{i}\in mcs(\mathrm{\Sigma})\mid\alpha\in\mathrm{\Sigma}_{i}\right\}\right|$
@@
Example(continue) Ontology $\mathrm{\Sigma}=\{\varphi_{1},\varphi_{2},\varphi_{3},\varphi_{4}\}$. $\mathrm{\Sigma}$ has 4 MCSs namely: $\{\varphi_{2},\varphi_{3},\varphi_{4}\}$, $\{\varphi_{1},\varphi_{3},\varphi_{4}\}$, $\{\varphi_{1},\varphi_{2},\varphi_{4}\}$, $\{\varphi_{1},\varphi_{2},\varphi_{3}\}$. We have #$\#mc(\mathrm{\Sigma},\varphi_{1})=\#mc(\mathrm{\Sigma},\varphi_{2})=\#mc(\mathrm{\Sigma},\varphi_{3})=\#mc(\mathrm{\Sigma},\varphi_{4})=3$, so the score of $\mathrm{\Sigma}_{1}(resp.\mathrm{\Sigma}_{2},\mathrm{\Sigma}_{3},\mathrm{\Sigma}_{4})$ w.r.t $score_{\mathrm{\Sigma},sum}^{\#mc}$ is equal to 9 (resp. 9, 9, 9). These four MCSs all have the same value and are selected for reasoning, i.e., an axiom can be inferred if it can be inferred from each of these four MCSs.
@@ 4.1. Semantic Representation & Embedding
Ontology Elements
OWL Representation Phrases or Sentences Triples
@@ 4.2. Semantic Similarity of Embedding
$Similarity$ represents certain similarity calculation method. The most common ones are based on Cosine Distance and Euclidean Distance:
$Similarity_{Cos}(\mathbf{v_{1}},\mathbf{v_{2}})=\frac{1}{2}(1+\frac{\mathbf{v_{1}}\cdot\mathbf{v_{2}}}{||\mathbf{v_{1}}||\times||\mathbf{v_{2}}||})$
$Similarity_{Euc}(\mathbf{v_{1}},\mathbf{v_{2}})=\frac{1}{1+\sqrt{||\mathbf{v_{1}}-\mathbf{v_{2}}||}}$
@@
Example (continue) By applying the cosine similarity metric and Sentence-BERT, we obtain $Sim_{Cos}(\varphi_{1},\varphi_{2})=0.41$ and $Sim_{Cos}(\varphi_{1},\varphi_{4})$ $=0.62$. This may be because $\varphi_{4}$ has a same concept ArtifactualFeatureType with $\varphi_{1}$ while $\varphi_{3}$ doesn’t, so $\varphi_{4}$ has higher semantic association with $\varphi_{1}$.
@@ 5. Logical Properties
Definition 9 ($score^{s}_{\Sigma,\oplus}$). Let $\mathrm{s}$ be a scoring function defined in Definition 2 and let $\oplus$ be an aggregation function. Let $\mathrm{\Sigma}$ be an ontology and   $\mathrm{\Sigma}_{i}=\{\alpha_{1}...\alpha_{n}\}\subseteq\mathrm{\Sigma}$.  We define $score^{s}_{\mathrm{\Sigma},\oplus}(\mathrm{\Sigma}_{i})=\oplus_{\alpha\in\mathrm{\Sigma}_{i}}s(\mathrm{\Sigma},\alpha)$.
@@
[…]
@@
[…]
@@
[…]
@@ 6. Experiment and Evaluation
AUTO.-cocus-eda $158$ $86$ $114$ $907$ $6$ $\mathcal{ALCOIN}$
@@ 6.1. Data
AUT.-cocus-eda is constructed in (Ji et al., 2014) by merging two source ontologies cocus and edas with their mapping generated by the mapping system AUTOMSv2 which has participated in the famous contest of ontology alignment evaluation initiative666 http://oaei.ontologymatching.org/2012/conference/. Its inconsistency is caused by multiple sources, lying on two similar concepts ‘cocus#person’ and ‘edas#person’. Bioportal-metadata777 https://bioportal.bioontology.org/ is a real-life ontology from the world’s most comprehensive repository of biomedical ontologies (Ji et al., 2014). Some inconsistencies come from multiple sources of certain individual. For instance, these four axioms are inconsistent:
@@
[…]

### v3

@@ 3. Preliminaries
Definition 2 (scoring function). A scoring function s associates with an ontology $\mathrm{\Sigma}$ and an axiom $\alpha\in\mathrm{\Sigma}$ a non-negative real number $s(\mathrm{\Sigma},\alpha)$ which is equal to 0 if and only if $\alpha$ is trivial (i.e., such that $\alpha$ is a tautology or a contradiction).
@@
$mc(\mathrm{\Sigma},\alpha)=\left|\left\{\mathrm{\Sigma}_{i}\in mcs(\mathrm{\Sigma})\mid\alpha\in\mathrm{\Sigma}_{i}\right\}\right|$
@@
Example(continue) Ontology $\mathrm{\Sigma}=\{\varphi_{1},\varphi_{2},\varphi_{3},\varphi_{4}\}$. $\mathrm{\Sigma}$ has 4 MCSs namely: $\{\varphi_{2},\varphi_{3},\varphi_{4}\}$, $\{\varphi_{1},\varphi_{3},\varphi_{4}\}$, $\{\varphi_{1},\varphi_{2},\varphi_{4}\}$, $\{\varphi_{1},\varphi_{2},\varphi_{3}\}$. We have $mc(\mathrm{\Sigma},\varphi_{1})=\#mc(\mathrm{\Sigma},\varphi_{2})=\#mc(\mathrm{\Sigma},\varphi_{3})=\#mc(\mathrm{\Sigma},\varphi_{4})=3$, so the score of $\mathrm{\Sigma}_{1}(resp.\mathrm{\Sigma}_{2},\mathrm{\Sigma}_{3},\mathrm{\Sigma}_{4})$ w.r.t $score_{\mathrm{\Sigma},sum}^{\#mc}$ is equal to 9 (resp. 9, 9, 9). These four MCSs all have the same value and are selected for reasoning, i.e., an axiom can be inferred if it can be inferred from each of these four MCSs.
@@ 4.1. Semantic Representation & Embedding
Ontology Elements OWL Representation Phrases or Sentences Triples
@@ 4.2. Semantic Similarity of Embedding
SimilaritySimilarity represents certain similarity calculation method. The most common ones are based on Cosine Distance and Euclidean Distance:
[…]
@@
Example (continue) By applying the cosine similarity metric and Sentence-BERT, we obtain $Sim_{Cos}(\varphi_{1},\varphi_{2})=0.41$ and $Sim_{Cos}(\varphi_{1},\varphi_{4})$ =0.62=0.62. This may be because $\varphi_{4}$ has a same concept ArtifactualFeatureType with $\varphi_{1}$ while $\varphi_{3}$ doesn’t, so $\varphi_{4}$ has higher semantic association with $\varphi_{1}$.
@@ 5. Logical Properties
Definition 9 ($score^{s}_{\Sigma,\oplus}$). Let $\mathrm{s}$ be a scoring function defined in Definition 2 and let $\oplus$ be an aggregation function. Let $\mathrm{\Sigma}$ be an ontology and $\mathrm{\Sigma}_{i}=\{\alpha_{1}...\alpha_{n}\}\subseteq\mathrm{\Sigma}\$. We define $score^{s}_{\mathrm{\Sigma},\oplus}(\mathrm{\Sigma}_{i})=\oplus_{\alpha\in\mathrm{\Sigma}_{i}}s(\mathrm{\Sigma},\alpha)$.
@@
Definition 12 ($mcs_{\succeq_{\mathrm{\Sigma}}}$). Given an ontology $\mathrm{\Sigma}$, an axiom $\alpha\in\mathrm{\Sigma}$, and a monotonic selection relation $\succeq_{\mathrm{\Sigma}}$, we define $mcs_{\succeq_{\mathrm{\Sigma}}}(\mathrm{\Sigma},\alpha)=\{\mathrm{\Sigma}_{i}\in mcs(\mathrm{\Sigma},\alpha)|\textit{there exists}\textit{no}\ \mathrm{\Sigma}^{\prime}_{i}\in mcs(\mathrm{\Sigma},\alpha)\ \textit{such that}\ \mathrm{\Sigma}^{\prime}_{i}\succ_{\mathrm{\Sigma}}\mathrm{\Sigma}_{i}\}$.
@@
Theorem 2. A relation $\mid\sim$ is rational if and only if there exists an ontology $\mathrm{\Sigma}$ and a monotonic selection relation $\succeq_{\Sigma}$ such that $\mid\sim_{\Sigma}^{{mcs}_{\succeq\Sigma}}\ =\ \mid\sim$.
@@
Theorem 3. The relation $\mid\sim$ defined by the selection relation which takes our proposed scoring function is rational.
@@ 6. Experiment and Evaluation
AUTO.-cocus-eda $158$ $86$ $114$ $907$ 66 $\mathcal{ALCOIN}$
@@ 6.1. Data
AUT.-cocus-eda is constructed in (Ji et al., 2014) by merging two source ontologies cocus and edas with their mapping generated by the mapping system AUTOMSv2 which has participated in the famous contest of ontology alignment evaluation initiative [^6] http://oaei.ontologymatching.org/2012/conference/. Its inconsistency is caused by multiple sources, lying on two similar concepts ‘cocus#person’ and ‘edas#person’. Bioportal-metadata [^7] https://bioportal.bioontology.org/ is a real-life ontology from the world’s most comprehensive repository of biomedical ontologies (Ji et al., 2014). Some inconsistencies come from multiple sources of certain individual. For instance, these four axioms are inconsistent:
@@
         documentationmax 1rdfs:Literal)documentation\ max\ 1\ rdfs:Literal)